OpenStudy (starryknight):

please help..... expalin in steps 4x2 - 32x + 16 = 0 answers; (x - 4)2 = 12 (x + 4)2 = -12 (x - 16)2 = 1 (x + 16)2 = -1

OpenStudy (anonymous):

Okay, we need to solve for x.

OpenStudy (anonymous):

okay i got you

OpenStudy (anonymous):

This is a quadratic equation, so we use the quadratic formula.

OpenStudy (anonymous):

I'll put it up here.

OpenStudy (anonymous):

\[x = \frac{ -b \sqrt{b^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

4x^2 - 32x + 16 = 0 b = 32 a = 4 c = 16

OpenStudy (anonymous):

So now we plug those values into the formula.

OpenStudy (anonymous):

\[x = \frac{ -32 \sqrt{b^{2}-4(4)(16)} }{ 2(4) }\]

OpenStudy (anonymous):

Oops, left out b for b^2 one second :D

OpenStudy (anonymous):

\[x = \frac{ -32 \sqrt{32^{2}-4(4)(16)} }{ 2(4) }\]

OpenStudy (anonymous):

So now we solve what's under the sqrt like so: \[\sqrt{32^{2} - 256 }\] \[\sqrt{1024 - 256}\] \[\sqrt{768}\]

OpenStudy (anonymous):

So now we have: \[x = \frac{ -32\sqrt{768} }{ 8 }\]

OpenStudy (starryknight):

thanx this was a long problem so thanx for sticking around :))

OpenStudy (anonymous):

No problem, almost done.

OpenStudy (starryknight):

kay ^^

OpenStudy (anonymous):

So now: \[x = \frac{ -512\sqrt{3} }{ }\]

OpenStudy (anonymous):

over 8

OpenStudy (anonymous):

x = -64 over sqrt 3

OpenStudy (phi):

notice in \[ 4x^2 - 32x + 16 = 0 \] that each term can be evenly divided by 4. so let's divide both sides of the equation by 4 (that means each term, on both sides) \[ \frac{4}{4} x^2 - \frac{32}{4} x + \frac{16}{4}= \frac{0}{4} \] which simplifies to \[ x^2 -8x +4 = 0\] based on the answer choices, it looks like we should COMPLETE THE SQUARE do you know how to do that ?

OpenStudy (anonymous):

I didn't even see the answer choices. *facepalm*

OpenStudy (phi):

to complete the square, start by "moving" the 4 to the right side of the equation. You do this by add -4 to both sides: \[ x^2 -8x +4 - 4 = 0 - 4 \\ x^2 -8x = -4 \]

OpenStudy (phi):

the next step is look at the number in front of the x. The coefficient is -8 divide the coefficient by 2, and then square the result: -8/2 = -4 square -4: -4* -4 = +16 add +16 to both sides: \[ x^2 -8x +16= -4 +16\] the left side is a perfect square. the right side simplifies